# A Survey on Mixture of Experts

### v1

@@ 4.1.3. Soft
Token Merging. Puigcerver et al. (Puigcerver et al., 2023) proposed the Soft MoE, which eschews the conventional sparse and discrete gating mechanism in favor of a soft assignment strategy that merges tokens. This method computes several weighted averages of all tokens, with weights depending on both tokens and experts, and processes each aggregate with its respective expert. Their experimental results in image classification demonstrate that soft MoE enhances the stability of gating function training and inherently maintains balance. HOMOE (Dat et al., 2023) follows the design of Soft MoE and combines it with Hopfield network to address the the challenges of Compositional Zero-Shot Learning tasks. Yet, merging input tokens complicates its application in auto-regressive decoders, as future tokens required for averaging are inaccessible during inference.
@@
Expert Merging. In contrast to the merging of input tokens, Muqeeth et al. (Muqeeth et al., 2023) introduced the Soft Merging of Experts with Adaptive Routing (SMEAR) framework, which circumvents discrete gating by merging all the experts’ parameters through a weighted average, as illustrated in Figure 4 (f). They argue that conventional sparse MoE models often fail to match the performance of their parameter-matched dense counterparts or those utilizing non-learned heuristic gating functions, potentially due to flawed gradient estimation methods for training modules with non-differentiable, discrete gating decisions. By processing the input tokens through a single merged expert, SMEAR does not incur a significant increase in computational costs and enables standard gradient-based training. Empirical evaluations on T5-GLUE and ResNet-DomainNet benchmarks reveal that SMEAR-equipped models surpass those with metadata-based (Gururangan et al., 2022; Kudugunta et al., 2021) or gradient-estimated learning gating strategies. On ResNet-DomainNet, SMEAR achieved a 1.5% higher average accuracy than Soft MoE (Puigcerver et al., 2023) with single “slot” per expert, at the expense of a near 10% reduction in throughput. Subsequent contributions by Zhong et al. (Zhong et al., 2024) argue that SMEAR’s demonstrated advantages are confined to downstream fine-tuning on classification tasks. They present Lory, an innovative approach for scaling such expert merging architectures to auto-regressive language model pre-training. Lory (Zhong et al., 2024) introduces a causal segment routing strategy, conducting expert merging at the segment level while maintaining the auto-regressive nature of language models. Furthermore, it employs similarity-based data batching to direct expert specialization in particular domains or topics. Lory’s empirical validation on LLaMA models showcases significant improvements over parameter-matched dense models in terms of perplexity (by 13.9%) and on diverse downstream tasks (by 1.5%-11.1%), highlighting the potential of fully-differentiable MoE architectures for language model pre-training and encouraging further investigation in this area. In addition, expert merging methods have demonstrated efficacy in parameter-efficient fine-tuning (PEFT) MoE contexts. Zadouri et al. (Zadouri et al., 2023) substantiate that soft merging of experts significantly outperforms sparse gating mechanisms (top-1, top-2) in the T5 models (Raffel et al., 2020) fine-tuning with the MoV-10 setting of 10 (IA)3 vector expert. Wu et al. (Wu et al., 2023a) propose Omni-SMoLA, an architecture leveraging the soft method to mix multimodal low-rank experts, improving the generalist performance across a broad range of generative vision-language tasks.
4.2. Experts
@@
4.2.1. Network Types
Since the initial integration of MoE into transformer architectures (Lepikhin et al., 2020; Fedus et al., 2022b; Zoph et al., 2022), MoE has served as a substitute for Feed-Forward Network (FFN) modules within these models. Typically, each expert within a MoE layer replicates the architecture of the FFN it replaces. This paradigm, wherein FFNs are utilized as experts, remains predominant, and subsequent refinements will be expounded upon in Sections 4.2.2 to 4.2.4.
Feed-Forward Network. As discussed in existing work (Sukhbaatar et al., 2024), the predilection for leveraging MoE in the context of FFNs is rooted in the hypothesis that self-attention layers exhibit lower sparsity and less domain specificity than FFN layers. Pan et al. (Pan et al., 2024) provide empirical support for this, revealing marked sparsity in FFN layers compared to self-attention layers, through their analysis of downstream Wikitext tasks using their pre-trained DS-MoE models. Their results indicate a mere 20% active expert engagement in FFN layers, in contrast to the 80% observed within self-attention layers. In earlier investigation of FFN computational patterns, Zhang et al. (Zhang et al., 2022a) observe that most inputs only activate a small proportion of neurons of FFNs, thus corroborating the inherent sparsity of FFNs.
Attention. While the focus of MoE research has predominantly been on FFN layers within the Transformer architecture, Zhang et al. (Zhang et al., 2022b) introduce the Mixture of Attention Heads (MoA), an innovative architecture that combines multi-head attention layers with MoE to further enhance performance and restrain computational cost. As delineated in Figure 5 (a), MoA employs two sets of experts, one for query projection and one for output projection, which are selected the same indices of experts through a common gating network. To reduce computational complexity, MoA shares the key () and value () projection weights across attention experts, with experts differentiated only by their respective query () and output () projection weights, allowing for shared pre-computation of key () and value () sequences. Subsequent work such as DS-MoE (Pan et al., 2024), JetMoE (Shen et al., 2024), and ModuleFormer (Shen et al., 2023c) follows the design of MoA and further refines the combination of MoE and attention layer.
Others. In addition to the aforementioned expert network types, researchers have explored the use of Convolutional Neural Network (CNN) as expert (Chowdhury et al., 2023; Zhang et al., 2023; Chen et al., 2022a; Wang et al., 2020; Gross et al., 2017). Moreover, recent endeavors that integrate Parameter-Efficient Fine-Tuning (PEFT) techniques with MoE, such as employing Low-Rank Adaptation (LoRA) (Hu et al., 2021) as expert, have shown promising results, which are discussed in Section 4.3.
@@ 4.2.1. Network Types
| Qwen1.5-MoE (Team, 2024b) (2024) | 2.7B/14.3B | 8/64 | 2048 | 5632 | 1408 | 24 | 16 | 128 | 1 | SwiGLU | 4 |
| DBRX (Databricks, 2024) (2024) | 36B/132B | 4/16 | 6144 | 10752 | 40 | 48 | 128 | 1 | SwiGLU | 0 | |
| Jamba (Lieber et al., 2024) (2024) | 12B/52B | 2/16 | 4096 | 14336 | 32 | 32 | 128 | 1/2, 1:7 Attention:Mamba | SwiGLU | 0 | |
| Skywork-MoE (Tianwen Wei, 2024) (2024) | 22B/146B | 2/16 | 4608 | 12288 | 52 | 36 | 128 | 1 | SwiGLU | 0 | |
| Yuan 2.0-M32 (Wu et al., 2024b) (2024) | 3.7B/40B | 2/32 | 2048 | 8192 | 24 | 16 | 256 | 1 | SwiGLU | 0 |
4.2.2. Hyperparameters
@@
The scale of sparse MoE models is governed by several critical hyperparameters that extend beyond those of dense transformer models. These include (1) the count of experts per MoE layer, (2) the size of each expert, and (3) the placement frequency of MoE layers throughout the model. The selection of these hyperparameters is crucial, as it profoundly influences model performance and computational efficiency across various tasks. Optimal hyperparameter choices are thus contingent upon the specific application requirements and the constraints of the computational infrastructure. Our subsequent analysis, informed by the exemplified models listed in Table 2, explores these hyperparameter decisions in depth. Meanwhile, we enumerate some recent open-source models, detailing their number of parameters and benchmark performance in Table 3.
@@
Expert Count. Initial investigations employing thousands of experts per layer yielded impressive gains in pre-training and translation quality (Shazeer et al., 2017; Lepikhin et al., 2020; Fedus et al., 2022b). Nonetheless, the quality of sparse MoE models is disproportionately reduced under domain shift (Artetxe et al., 2021) or when fine-tuning on diverse task distributions (Fedus et al., 2022b). GLaM (Du et al., 2022) adopts a configuration of 64 experts, guided by their findings that a 64-expert setup with top-2 gating strikes an optimal balance between execution efficiency and performance across zero-shot, one-shot, and few-shot scenarios. Reflecting this trend, more recent sparse MoE models (Zoph et al., 2022; Jiang et al., 2024a; Xue et al., 2024; Team, 2024b; Databricks, 2024; Tianwen Wei, 2024; Lieber et al., 2024; Wu et al., 2024b) commonly utilize no more than 64 experts. Additionally, DeepSpeed-MoE (Rajbhandari et al., 2022) adopts a Pyramid-MoE approach, positioning MoE layers with a larger expert count towards the network’s end.
@@
Expert Size. To scale the model effectively, GLaM (Du et al., 2022) prioritizes the expansion of the intermediate hidden dimension per expert while standardizing the expert count at 64, a strategy that often requires the implementation of tensor parallelism across multiple accelerators to maintain computational efficiency (Du et al., 2022; Rajbhandari et al., 2022; Fedus et al., 2022b). From this period forward, MoE models (Zoph et al., 2022; Jiang et al., 2024a; Databricks, 2024; Tianwen Wei, 2024) typically featured larger expert dimensions. Differently, DeepSeekMoE (Dai et al., 2024; DeepSeek-AI, 2024) introduces the concept of fine-grained expert segmentation by subdividing the intermediate hidden dimension of FFN expert, while preserving the overall parameter count. Specifically, DeepSeekMoE-145B employs a reduced intermediate hidden dimension at one-eighth that of its dense FFN counterpart, increasing both the number of experts (from 16 to 128) and the number of active experts (from top-2 to top-16) by a factor of eight. They believe that this strategy not only refines the decomposition of knowledge across experts, facilitating more precise learning, but also enhances the flexibility of expert activation combinations, allowing for more specialized and targeted knowledge capture. Qwen1.5-MoE (Team, 2024b) and DBRX (Databricks, 2024) adopt a similar fine-grained expert segmentation strategy. From the results of LLAMA-MoE, which evenly splits the parameters of FFNs into non-overlapping experts to construct the MoE models with same parameter count, activating 4 of 16 expert performs slightly better than activating 2 of 8 expert. Results from LLAMA-MoE (Team, 2023), which allocates dense FFN parameters across non-overlapping experts to maintain a consistent parameter count, indicate that activating 4 out of 16 experts with a dimensionality of marginally outperforms the activation of 2 out of 8 experts with .
@@
Frequency of MoE Layers. Sparse MoE models typically evolve from dense architectures by interspersing MoE layers in place of the dense FFN layers at regular intervals. Although a higher frequency of MoE layers can enlarge the model size, it also introduces greater system overhead. In practice, most MoE models features alternate FFN replacement (1/2) with MoE layers (Lepikhin et al., 2020; Du et al., 2022; Artetxe et al., 2021; Rajbhandari et al., 2022). Nevertheless, variations exist, with some models incorporating MoE layers every fourth layer (1/4) (Zoph et al., 2022; Xue et al., 2024) or in every layer (1/1) (Fedus et al., 2022b; Dai et al., 2024). Following the introduction of Mixtral 8x7B (Jiang et al., 2024a), the trend seems to shift towards placing MoE in every layer of the model, with a common choice of only 8 or 16 experts mirroring the dimensionality of a dense FFN (Dai et al., 2024; Team, 2024b; Databricks, 2024; Tianwen Wei, 2024).
@@
Research into the optimal configuration of MoE layers has been extensive. V-MoE (Riquelme et al., 2021) employs MoE in the last few even-numbered Transformer layers, noting that, despite using fewer MoE layers, the impact on performance is minimal while computational speed is significantly enhanced. DeepSeekMoE-16B/-145B (Dai et al., 2024) replaces all FFNs with MoE layers, excluding the first, due to the observed slower convergence of load balance status in the first layer. MoE-LLaVA (Lin et al., 2024), a recently popular open Large Vision-Language Model (LVLM), demonstrates that alternating MoE placement yields superior model quality and execution efficiency on multimodal tasks, compared to every-layer MoE placement or ”First-Half” and ”Second-Half” configurations. ST-MoE (Zoph et al., 2022) found that adding a dense FFN adjacent to each MoE layer can improve model quality. Brainformers (Zhou et al., 2023) introduce a nonuniform architecture that integrates MoE layers, dense FFNs, attention mechanisms, and a variety of layer normalizations and activation functions without strict sequential layering, trading architectural regularity for the flexibility of sub-layer composition. Jamba (Lieber et al., 2024) integrates the architecture of Mamba (Gu and Dao, 2023) by adopting a 1:7 ratio of attention-to-Mamba layers.
@@ 4.2.3. Activation Function
Building upon dense Transformer architectures, sparse MoE models have adopted a progression of activation functions paralleling those in leading dense large language models, including BERT (Devlin et al., 2018), T5 (Raffel et al., 2020), GPT (Brown et al., 2020), LLAMA (Touvron et al., 2023) and so on. The evolution of activation functions has seen a shift from ReLU (Glorot et al., 2011) to more advanced options such as GeLU (Hendrycks and Gimpel, 2016), GeGLU (Shazeer, 2020), and SwiGLU (Shazeer, 2020). This trend extends to other components of MoE models, which now frequently incorporate Root Mean Square Layer Normalization (RMSNorm) (Zhang and Sennrich, 2019), Grouped Query Attention (GQA) (Ainslie et al., 2023), and Rotary Position Embeddings (RoPE) (Su et al., 2024).
4.2.4. Shared Expert

### v2

@@ 4.1.3 Soft
Token Merging. Puigcerver et al. [39] proposed the Soft MoE, which eschews the conventional sparse and discrete gating mechanism in favor of a soft assignment strategy that merges tokens. This method computes several weighted averages of all tokens, with weights depending on both tokens and experts, and processes each aggregate with its respective expert. Their experimental results in image classification demonstrate that soft MoE enhances the stability of gating function training and inherently maintains balance. HOMOE [96] follows the design of Soft MoE and combines it with Hopfield network to address the the challenges of Compositional Zero-Shot Learning tasks. Yet, merging input tokens complicates its application in auto-regressive decoders, as future tokens required for averaging are inaccessible during inference.
@@
Expert Merging. In contrast to the merging of input tokens, Muqeeth et al. [40] introduced the Soft Merging of Experts with Adaptive Routing (SMEAR) framework, which circumvents discrete gating by merging all the experts’ parameters through a weighted average, as illustrated in Figure 4 (f). They argue that conventional sparse MoE models often fail to match the performance of their parameter-matched dense counterparts or those utilizing non-learned heuristic gating functions, potentially due to flawed gradient estimation methods for training modules with non-differentiable, discrete gating decisions. By processing the input tokens through a single merged expert, SMEAR does not incur a significant increase in computational costs and enables standard gradient-based training. Empirical evaluations on T5-GLUE and ResNet-DomainNet benchmarks reveal that SMEAR-equipped models surpass those with metadata-based [91, 77] or gradient-estimated learning gating strategies. On ResNet-DomainNet, SMEAR achieved a 1.5% higher average accuracy than Soft MoE [39] with single “slot” per expert, at the expense of a near 10% reduction in throughput. Subsequent contributions by Zhong et al. [41] argue that SMEAR’s demonstrated advantages are confined to downstream fine-tuning on classification tasks. They present Lory, an innovative approach for scaling such expert merging architectures to auto-regressive language model pretraining. Lory [41] introduces a causal segment routing strategy, conducting expert merging at the segment level while maintaining the auto-regressive nature of language models. Furthermore, it employs similarity-based data batching to direct expert specialization in particular domains or topics. Lory’s empirical validation on LLaMA models showcases significant improvements over parameter-matched dense models in terms of perplexity (by 13.9%) and on diverse downstream tasks (by 1.5%-11.1%), highlighting the potential of fully-differentiable MoE architectures for language model pretraining and encouraging further investigation in this area. In addition, expert merging methods have demonstrated efficacy in parameter-efficient fine-tuning (PEFT) MoE contexts. Zadouri et al. [42] substantiate that soft merging of experts significantly outperforms sparse gating mechanisms (top-1, top-2) in the T5 models [171] fine-tuning with the MoV-10 setting of 10 (IA)3 vector expert. Wu et al. [43] propose Omni-SMoLA, an architecture leveraging the soft method to mix multimodal low-rank experts, improving the generalist performance across a broad range of generative vision-language tasks. He et al. [97] introduce Merging Experts into One (MEO), merging multiple selected experts into one to reduce the expert computation cost. Moreover, they perform the expert selection at the sequence level and employ a token attention mechanism for capturing the identification of each token, thus preserving context information without the necessity of merging distinct weights and biases for individual tokens.
4.2 Experts
@@
4.2.1 Network Types
Since the initial integration of MoE into transformer architectures [28, 36, 37], MoE has served as a substitute for Feed-Forward Network (FFN) modules within these models. Typically, each expert within a MoE layer replicates the architecture of the FFN it replaces. This paradigm, wherein FFNs are utilized as experts, remains predominant, and subsequent refinements will be expounded upon in Sections 4.2.2 to 4.2.4.
Feed-Forward Network. As discussed in existing work [54], the predilection for leveraging MoE in the context of FFNs is rooted in the hypothesis that self-attention layers exhibit lower sparsity and less domain specificity than FFN layers. Pan et al. [64] provide empirical support for this, revealing marked sparsity in FFN layers compared to self-attention layers, through their analysis of downstream Wikitext tasks using their pretrained DS-MoE models. Their results indicate a mere 20% active expert engagement in FFN layers, in contrast to the 80% observed within self-attention layers. In earlier investigation of FFN computational patterns, Zhang et al. [50] and Li et al. [172] observe that most inputs only activate a small proportion of neurons of FFNs, highlighting the inherent sparsity of FFNs. Subsequently, Zhang et al. [98] observe the Emergent Modularity (EM) phenomenon within pretrained Transformers, revealing a significant correlation between neuron activation and specific tasks (evidenced by the functional specialization of neurons and function-based neuron grouping). This discovery supports the proposition that the MoE structure reflects the modularity of pre-trained Transformers.
Attention. While the focus of MoE research has predominantly been on FFN layers within the Transformer architecture, Zhang et al. [82] introduce the Mixture of Attention Heads (MoA), an innovative architecture that combines multi-head attention layers with MoE to further enhance performance and restrain computational cost. As delineated in Figure 5 (a), MoA employs two sets of experts, one for query projection and one for output projection. Both sets select the experts with the same indices through a common gating network. To reduce computational complexity, MoA shares the key () and value () projection weights across attention experts, with experts differentiated only by their respective query () and output () projection weights, allowing for shared pre-computation of key () and value () sequences. Subsequent work such as DS-MoE [64], JetMoE [83], and ModuleFormer [73] follows the design of MoA and further refines the combination of MoE and attention layer.
Others. In addition to the aforementioned expert network types, researchers have explored the use of Convolutional Neural Network (CNN) as expert [99, 100, 101, 102, 173]. Moreover, recent endeavors that integrate Parameter-Efficient Fine-Tuning (PEFT) techniques with MoE, such as employing Low-Rank Adaptation (LoRA) [174] as expert, have shown promising results, which are discussed in Section 4.3.
@@ 4.2.1 Network Types
Qwen1.5-MoE [104] (2024) | 2.7B/14.3B | 8/64 | 2048 | 5632 | 1408 | 24 | 16 | 128 | 1 | SwiGLU | 4 |
DBRX [31] (2024) | 36B/132B | 4/16 | 6144 | 10752 | 40 | 48 | 128 | 1 | SwiGLU | 0 | |
Jamba [68] (2024) | 12B/52B | 2/16 | 4096 | 14336 | 32 | 32 | 128 | 1/2, 1:7 Attention:Mamba | SwiGLU | 0 | |
Skywork-MoE [67] (2024) | 22B/146B | 2/16 | 4608 | 12288 | 52 | 36 | 128 | 1 | SwiGLU | 0 | |
Yuan 2.0-M32 [84] (2024) | 3.7B/40B | 2/32 | 2048 | 8192 | 24 | 16 | 256 | 1 | SwiGLU | 0 |
4.2.2 Hyperparameters
@@
The scale of sparse MoE models is governed by several critical hyperparameters that extend beyond those of dense transformer models. These include (1) the count of experts per MoE layer, (2) the size of each expert, and (3) the placement frequency of MoE layers throughout the model. The selection of these hyperparameters is crucial, as it profoundly influences model performance and computational efficiency across various tasks. Optimal hyperparameter choices are thus contingent upon the specific application requirements and the constraints of the computational infrastructure. Our subsequent analysis, informed by the exemplified models listed in Table II, explores these hyperparameter decisions in depth. Meanwhile, we enumerate some recent open-source models, summarizing their number of parameters and benchmark performance in Table III.
@@
Expert Count. Initial investigations employing thousands of experts per layer yielded impressive gains in pretraining and translation quality [27, 28, 36]. Nonetheless, the quality of sparse MoE models is disproportionately reduced under domain shift [103] or when fine-tuning on diverse task distributions [36]. GLaM [35] adopts a configuration of 64 experts, guided by their findings that a 64-expert setup with top-2 gating strikes an optimal balance between execution efficiency and performance across zero-shot, one-shot, and few-shot scenarios. Reflecting this trend, more recent sparse MoE models [37, 29, 38, 104, 31, 67, 68, 84] commonly utilize no more than 64 experts. Additionally, DeepSpeed-MoE [66] adopts a Pyramid-MoE approach, positioning MoE layers with a larger expert count towards the network’s end.
@@
Expert Size. To scale the model effectively, GLaM [35] prioritizes the expansion of the intermediate hidden dimension per expert while standardizing the expert count at 64, a strategy that often requires the implementation of tensor parallelism across multiple accelerators to maintain computational efficiency [35, 66, 36]. From this period forward, MoE models [37, 29, 31, 67] typically featured larger expert dimensions. Differently, DeepSeekMoE [69, 33] introduces the concept of fine-grained expert segmentation by subdividing the intermediate hidden dimension of FFN expert, while preserving the overall parameter count. Specifically, DeepSeekMoE-145B employs a reduced intermediate hidden dimension at one-eighth that of its dense FFN counterpart, increasing both the number of experts (from 16 to 128) and the number of active experts (from top-2 to top-16) by a factor of eight. They believe that this strategy not only refines the decomposition of knowledge across experts, facilitating more precise learning, but also enhances the flexibility of expert activation combinations, allowing for more specialized and targeted knowledge capture. Qwen1.5-MoE [104] and DBRX [31] adopt a similar fine-grained expert segmentation strategy. Results from LLAMA-MoE [51], which allocates dense FFN parameters across non-overlapping experts to maintain a consistent parameter count, indicate that activating 4 out of 16 experts with a dimensionality of marginally outperforms the activation of 2 out of 8 experts with . Furthermore, Parameter Efficient Expert Retrieval (PEER) [105], an innovative layer design employing the product key technique [175] for sparse retrieval from a vast pool of tiny experts (over a million single-neuron experts), surpasses dense FFNs and coarse-grained MoEs in terms of performance-compute trade-off on language modeling tasks.
@@
Frequency of MoE Layers. Sparse MoE models typically evolve from dense architectures by interspersing MoE layers in place of the dense FFN layers at regular intervals. Although a higher frequency of MoE layers can enlarge the model size, it also introduces greater system overhead. In practice, most MoE models features alternate FFN replacement (1/2) with MoE layers [28, 35, 103, 66]. Nevertheless, variations exist, with some models incorporating MoE layers every fourth layer (1/4) [37, 38] or in every layer (1/1) [36, 69]. Following the introduction of Mixtral 8x7B [29], the trend seems to shift towards placing MoE in every layer of the model, with a common choice of only 8 or 16 experts mirroring the dimensionality of a dense FFN [69, 104, 31, 67].
@@
Research into the optimal configuration of MoE layers has been extensive. V-MoE [8] employs MoE in the last few even-numbered Transformer layers, noting that, despite using fewer MoE layers, the impact on performance is minimal while computational speed is significantly enhanced. DeepSeekMoE-16B/-145B [69] replaces all FFNs with MoE layers, excluding the first, due to the observed slower convergence of load balance status in the first layer. MoE-LLaVA [106], a recently popular open Large Vision-Language Model (LVLM), demonstrates that alternating MoE placement yields superior model quality and execution efficiency on multimodal tasks, compared to every-layer MoE placement or ”First-Half” and ”Second-Half” configurations. ST-MoE [37] found that adding a dense FFN adjacent to each MoE layer can improve model quality. Brainformers [95] introduce a nonuniform architecture that integrates MoE layers, dense FFNs, attention mechanisms, and a variety of layer normalizations and activation functions without strict sequential layering, trading architectural regularity for the flexibility of sub-layer composition. Jamba [68] integrates the architecture of Mamba [176] by adopting a 1:7 ratio of attention-to-Mamba layers.
@@ 4.2.3 Activation Function
Building upon dense Transformer architectures, sparse MoE models have adopted a progression of activation functions paralleling those in leading dense large language models, including BERT [181], T5 [171], GPT [2], LLAMA [161] and so on. The evolution of activation functions has seen a shift from ReLU [107] to more advanced options such as GeLU [108], GeGLU [109], and SwiGLU [109]. This trend extends to other components of MoE models, which now frequently incorporate Root Mean Square Layer Normalization (RMSNorm) [182], Grouped Query Attention (GQA) [183], and Rotary Position Embeddings (RoPE) [184].
4.2.4 Shared Expert